# Large Pre-trained time series models for cross-domain Time series analysis tasks

### v2

@@ Abstract
Large pre-trained models have been vital in recent advancements in domains like language and vision, making model training for individual downstream tasks more efficient and provide superior performance. However, tackling time-series analysis tasks usually involves designing and training a separate model from scratch leveraging training data and domain expertise specific to the task. We tackle a significant challenge for pre-training a foundational time-series model from multi-domain time-series datasets: extracting semantically useful tokenized inputs to the model across heterogenous time-series from different domains. We propose Large Pre-trained Time-series Models (LPTM) that introduces a novel method of adaptive segmentation that automatically identifies optimal dataset-specific segmentation strategy during pre-training. This enables LPTM to perform similar to or better than domain-specific state-of-art model when fine-tuned to different downstream time-series analysis tasks and under zero-shot settings. LPTM achieves superior forecasting and time-series classification results taking up to 40% less data and 50% less training time compared to state-of-art baselines.
@@ 1 Introduction
Time-series analysis tasks are important well-studied problems such as forecasting (Hyndman and Athanasopoulos, 2018) and classification (Chowdhury et al., 2022) with applications in wide-ranging domains such as retail, meteorology, economics, and health. Recent works (Chen et al., 2021; Wang et al., 2022; Zeng et al., 2023) have shown the efficacy of purely data-driven deep learning models. However, most state-of-art neural models are known to be data-hungry and require substantial training data. Motivated by language and vision foundational models Bommasani et al. (2021), recent body of works build pre-trained time-series models Das et al. (2023); Ansari et al. (2024); Jin et al. (2024); Rasul et al. (2023). These models are trained on diverse datasets from different domains during pre-training. They require less training resources and data and produce superior performance. These models can also be deployed without any training, in a zero-shot or few-shot setting.
@@
Most neural sequential models input time-series values for each time-step separately. However, unlike text data, each individual time stamp may not provide enough semantic meaning about local temporal patterns of the time series. To tackle this, Nie et al. (2022) proposed to segment the time series and input each segment as individual tokens to their transformer-based model. This simple segmentation method of tokenizing time-series has been used by recent pre-trained models Woo et al. (2024); Das et al. (2023) to provide superior performance across multiple applications. However, segmenting input time-series uniformly with fixed-length segments, while simple, can be a very inflexible tokenizing method, especially when dealing with datasets from multiple domains with different set of underlying generative dynamics, sampling rate, noise, etc. For example, among two datasets, a dataset with a lower sampling rate (such as GDP time-series) may require longer segments than those with higher sampling rates to capture similar patterns in the model (such as heart sensors collecting data in milliseconds). However, note that the dynamics of the same time-series may vary with time. For, time intervals that are smoother with less complex dynamics, using longer segment sizes may suffice but intervals where time-series are complex and have multiple temporal patterns may require finer-grained segmentation. For example, for seasonal epidemics, time-series is smoother during the off-season and has more complex dynamics during outbreaks and times of higher incidence.
@@ 4 Related Works
Recent works have shown the efficacy of self-supervised representation learning for time-series for various classification and forecasting tasks in a wide range of applications such as modeling behavioral datasets Merrill and Althoff (2022); Chowdhury et al. (2022), power generation Zhang et al. (2019), health care Zhang et al. (2022). Franceschi et al. (2019) used triplet loss to discriminate segments of the same time-series from others. These works use methods such as contrastive losses Eldele et al. (2021); Yue et al. (2022) or other similarity metric-based techniques Tonekaboni et al. (2021). However, all these methods apply SSL on the same dataset that is used for training and may not adapt well to multi-domain datasets. There has been some recent works leveraging foundational models like LLMs for time-series forecasting across multiple applications. One set of works directly uses LLMs without fine-tuning to perform time-series forecasting via careful prompting Gruver et al. (2024); Jin et al. (2024). Other works fine-tune LLMs specifically for time-series forecasting Zhou et al. (2023); Rasul et al. (2023); Ansari et al. (2024). The time-series representation used by these models includes using individual time-steps as input Ansari et al. (2024), converting each digit of the time-series to character embeddings to be directly used by LLMs Gruver et al. (2024); Jin et al. (2024) and uniform segmentation Das et al. (2023); Woo et al. (2024); Zhou et al. (2023). LPTM performs superior to these methods while being 10x to 100x smaller than the large LLMs used as backbones.
@@ 5 Experiment Setup
We derive pre-train time-series datasets from multiple domains: Epidemics: We use a large number of epidemic time-series aggregated by Project Tycho (van Panhuis et al., 2018). from 1888 to 2021 for different diseases collected at state and city levels in the US. We remove time series with missing data and use time series for 11 diseases of very diverse epidemic dynamics such as seasonality, biology, geography, etc.: Hepatitis A, measles, mumps, pertussis, polio, rubella, smallpox, diphtheria, influenza, typhoid and Cryptosporidiosis (Crypto.). Electricity: We use ETT electricity datasets (ETT1 and ETT2) collected from (Zhou et al., 2021) at 1 hour intervals over 2 years. We use the default 12/4/4 train/val/test split and use the train split for pre-training as well. Traffic Datasets: We use 2 datasets related to traffic speed prediction. PEMS-Bays and METR-LA (Li et al., 2017) are datasets of traffic speed at various spots collected by the Los Angeles Metropolitan Transportation Authority and California Transportation Agencies over 4-5 months. Demand Datasets: We use bike and taxi demand datasets from New York City collected from April to June 2016 sampled every 30 minutes. We all but the last 5 days of data for training and pre-training. Stock forecasting: We also collect the time-series of daily stock prices of Nasdaq and S&P 500 index using Yahoo finance data (Xu and Berkely, 2014) from July 2014 to June 2019. We train and pre-train using the first 800 trading days and use the last 400 for testing. M4 competition time-series: We also used the 3003 time-series of M4 forecasting competition (Makridakis and Hibon, 2000) which contains time-series from multiple domains including demographics, finance, and macroeconomics. Motion and behavioral sensor datasets: We use the set of sensor datasets extracted from UEA archive (Bagnall et al., 2018) and UCI Machine learning repository (Asuncion and Newman, 2007) similar to (Chowdhury et al., 2022). Note that our publicly accessible pre-training dataset is significantly smaller than other pre-trained datasets used by past work Ansari et al. (2024); Das et al. (2023) some of which use confidential data inaccessible to us. We also do not use any synthetic datasets like Das et al. (2023); Ansari et al. (2024).
@@
We compared LPTM’s performance in various time-series tasks against eighteen state-of-the-art general forecasting and domain-specific baselines. First, we compare against recent pre-trained foundational time-series models: (1) LLM-Time Gruver et al. (2024), (2) TimesFM Das et al. (2023), (3) Lag-LLAMA Rasul et al. (2023) and (4) Chronos Ansari et al. (2024). We skip models like Time-LLM Jin et al. (2024) and Autotimes Liu et al. (2024) which cannot perform zero-shot forecasting across domains and are outperformed by the aforementioned recent models when fine-tuned. We compared with (5) Informer Zhou et al. (2021), (6) Autoformer Chen et al. (2021), (7) PatchTST Nie et al. (2022), three state-of-the-art transformer-based forecasting models. We also compare against other recent model (8) MICN (Wang et al., 2022) and (9) TimesNeT Wu et al. (2023). We also compare with (10) N-HITS Challu et al. (2023) which uses multi-scale interpolation and (11) AutoARIMA Hyndman and Khandakar (2008) a ARIMA based model that does automatic hyperparameter search. We also compare it against three other state-of-art self-supervised methods for time-series: (12) TS2Vec (Yue et al., 2022), (13) TS-TCC (Eldele et al., 2021) and (14) SimMTM Dong et al. (2024) uses masking as pre-trained task for time-series classification.
@@ 6 Results
| AutoARIMA | 2.14 | 1344 | 0.73 | 0.64 | 4.1 | 4.13 | 16.43 | 0.62 | 1.89 |
| Informer | 1.62 | 1139 | 0.57 | 0.71 | 3.1 | 2.89 | 12.33 | 0.83 | 1.055 |
| Autoformer | 1.41 | 1227 | 0.72 | 0.82 | 2.7 | 2.73 | 12.71 | 0.19 | 0.887 |
@@
| TimesNet | 1.04 | 1194 | 0.56 | 0.62 | 3.9 | 2.83 | 11.82 | 0.19 | 1.055 |
| LLM-Time | 1.21 | 1319 | 0.52 | 0.49 | 3.9 | 3.7 | 12.11 | 0.21 | 1.064 |
| TimesFM | 1.32 | 1214 | 0.58 | 0.49 | 3.7 | 2.8 | 12.19 | 0.22 | 1.07 |
| Lag-LLAMA | 1.46 | 1416 | 0.61 | 0.57 | 3.9 | 2.9 | 13.43 | 0.28 | 1.33 |
| Chronos | 1.21 | 1228 | 0.59 | 0.52 | 3.7 | 3.1 | 12.82 | 0.27 | 1.04 |
@@
| EpiFNP | 0.52 | 872 | 0.81 | 1.25 | 4.1 | 2.98 | 12.11 | 0.28 | 1.281 |
| ColaGNN | 1.65 | 694 | 0.72 | 1.19 | 3.9 | 3.19 | 14.97 | 0.25 | 1.185 |
| TS2Vec | 1.85 | 905.9 | 0.99 | 1.74 | 3.5 | 3.11 | 13.48 | 0.94 | 1.344 |
| SimMTM | 1.31 | 1289 | 0.61 | 0.55 | 3.4 | 3.1 | 12.79 | 0.28 | 1.284 |
| TS-TCC | 1.94 | 1134.6 | 0.75 | 1.29 | 3.3 | 2.97 | 15.55 | 0.76 | 1.274 |
@@
| LPTM-NoSegment | 0.93 | 766 | 0.57 | 0.55 | 3.2 | 3.17 | 14.96 | 0.27 | 1.146 |
| LPTM-NoPreTrain | 0.96 | 827 | 0.46 | 0.57 | 3.7 | 2.66 | 12.43 | 0.25 | 1.271 |
| LPTM-NoLinProb | 0.92 | 885 | 0.43 | 0.53 | 3.1 | 2.49 | 12.17 | 0.19 | 1.032 |
| LPTM-OnlyRandMask | 0.87 | 895 | 0.51 | 0.52 | 2.8 | 2.42 | 12.36 | 0.21 | 1.076 |
| LPTM-OnlyLastMask | 0.79 | 773 | 0.44 | 0.48 | 2.7 | 2.31 | 12.04 | 0.19 | 0.916 |
The code for implementation of LPTM and datasets are provided at link333https://github.com/kage08/SegmentTS/ and hyperparameters are discussed in the Appendix. LPTM consists of 10 layers for the transformer and overall has about 100M parameters which 2x to over 10x smaller than other pre-trained time-series models.
@@
We also train the models with training data to evaluate fine-tuned performance. The forecasting performance are shown in Table 2. We evaluated LPTM against seventeen other forecasting baselines. LPTM is either the first or a close second best-performing model in all the benchmarks. LPTM generally outperforms general time-series forecasting methods as well as 2recent pre-trained models in all benchmarks despite its much lower parameter count (2-10x lower). Further, it is competitive or superior in performance to domain-specific methods designed specifically for the given domains (such as EpiFNP and STEP). LPTM beats the previous state-of-art domain-specific baselines in five of the benchmarks and comes second in four more. Finally, LPTM improves upon the state-of-art on electricity forecasting, traffic forecasting, and M3 datasets.
@@
Unlike other autoregressive foundational models designed for forecasting, LPTM can be used for classification due to its encoder-style architecture. We add a single classification layer over pooled output embeddings to predict the class logits. We evaluate LPTM and baselines on the classification of 35 sensor and behavioral datasets from UCI classification repositiry(Asuncion and Newman, 2007). We report the accuracy scores in Appendix Table 4. We observe that LPTM has highest mean rank and largest number of times it outperforms all baselines.
Data efficiency
Training efficiency
Ablation and Sensitivity Studies
Segments generated by LPTM
We also visualized the segments in Appendix Fig. 4. We observe that the segment sizes are smaller in regions of high variance or important parts of the time-series such as peak of the epidemic whereas simpler trends have longer segments which matches our intuition.
7 Conclusion
Our work mainly focuses on the important challenge of providing semantically meaningful inputs to the model that caters to learning time-series segmentation strategies specific to each domain. This is crucial when pre-training on diverse datasets, a key challenge for time-series data. The underlying model architecture is a straightforward transformer encoder that uses well-known masking techniques for self-supervised pre-training. Therefore, our method can be extended to leverage novel time-series model architectures and SSL methods. Extending our methods to provide calibrated forecasts that provide reliable uncertainty measures is also another important direction of research.
Appendix A Hyperparameters
Appendix B Data efficiency
| Informer | 27.3 | 25.5 | 45.1 | 49.7 | 27.1 | 49.6 | 17.5 | 14.3 |
| Autoformer | 19.5 | 29.3 | 49.5 | 55.2 | 18.5 | 45.1 | 11.9 | 19.7 |
| N-HITS | 15.4 | 22.5 | 36.1 | 49.3 | 26.4 | 47.5 | NA | NA |
| PatchTST | 12.9 | 29.5 | 36.4 | 45.7 | 18.2 | 49.4 | NA | NAg |
| MICN | 17.6 | 15.7 | 39.7 | 41.1 | 19.2 | 33.9 | NA | NA |
| TimesNet | 15.4 | 19.7 | 37.4 | 46.3 | 24.1 | 36.5 | 9.4 | 11.5 |
| STEP | 25.4 | 34.1 | 52.7 | 74.3 | 29.7 | 52.8 | NA | NA |
| EpiFNP | 22.5 | 39.5 | 41.1 | 39.1 | 21.6 | 97.6 | NA | NA |
| ColaGNN | 34.7 | 33.6 | 53.1 | 47.6 | 32.1 | 72.2 | NA | NA |
| TARNet | NA | NA | NA | NA | NA | NA | 13.7 | 9.4 |
| TS2Vec | 29.3 | 28.2 | 41.9 | 41.9 | 29.8 | 67.4 | 9.3 | 13.2 |
| TS-TCC | 21.7 | 23.7 | 46.3 | 44.3 | 25.3 | 55.8 | 12.7 | 11.1 |
| LPTM | 12.2 | 19.3 | 41.9 | 37.5 | 17.3 | 31.2 | 6.1 | 12.7 |
| LPTM-TB | NA | 12.5 | 29.6 | 32.9 | NA | 23.7 | 6.1 | 8.1 |
Appendix C Time-series Classification
| Informer | Autoformer | TimesNet | TARNet(SOTA) | TS2Vec | TS-TCC | TST | SimMTM | CRT | LPTM | |

### v3

@@ Abstract
Large pre-trained models have been vital in recent advancements in domains like language and vision, making model training for individual downstream tasks more efficient and provide superior performance. However, tackling time-series analysis tasks usually involves designing and training a separate model from scratch leveraging training data and domain expertise specific to the task. We tackle a significant challenge for pre-training a foundational time-series model from multi-domain time-series datasets: extracting semantically useful tokenized inputs to the model across heterogenous time-series from different domains. We propose Large Pre-trained Time-series Models (LPTM) that introduces a novel method of adaptive segmentation that automatically identifies optimal dataset-specific segmentation strategy during pre-training. This enables LPTM to perform similar to or better than domain-specific state-of-art model when fine-tuned to different downstream time-series analysis tasks and under zero-shot settings. LPTM achieves superior forecasting and time-series classification results taking up to 40% less data and 50% less training time compared to state-of-art baselines. Code: www.github.com/AdityaLab/Samay
@@ 1 Introduction
Time-series analysis tasks are important well-studied problems such as forecasting (Hyndman and Athanasopoulos, 2018) and classification (Chowdhury et al., 2022) with applications in wide-ranging domains such as retail, meteorology, economics, and health. Recent works (Chen et al., 2021; Wang et al., 2022; Zeng et al., 2023) have shown the efficacy of purely data-driven deep learning models. However, most state-of-art neural models are known to be data-hungry and require substantial training data. Motivated by language and vision foundational models Bommasani et al. (2021), recent body of works build pre-trained time-series models Das et al. (2023b); Ansari et al. (2024); Jin et al. (2024); Rasul et al. (2023). These models are trained on diverse datasets from different domains during pre-training. They require less training resources and data and produce superior performance. These models can also be deployed without any training, in a zero-shot or few-shot setting.
@@
Most neural sequential models input time-series values for each time-step separately. However, unlike text data, each individual time stamp may not provide enough semantic meaning about local temporal patterns of the time series. To tackle this, Nie et al. (2022) proposed to segment the time series and input each segment as individual tokens to their transformer-based model. This simple segmentation method of tokenizing time-series has been used by recent pre-trained models Woo et al. (2024); Das et al. (2023b) to provide superior performance across multiple applications. However, segmenting input time-series uniformly with fixed-length segments, while simple, can be a very inflexible tokenizing method, especially when dealing with datasets from multiple domains with different set of underlying generative dynamics, sampling rate, noise, etc. For example, among two datasets, a dataset with a lower sampling rate (such as GDP time-series) may require longer segments than those with higher sampling rates to capture similar patterns in the model (such as heart sensors collecting data in milliseconds). However, note that the dynamics of the same time-series may vary with time Liu et al. (2024a). For, time intervals that are smoother with less complex dynamics, using longer segment sizes may suffice but intervals where time-series are complex and have multiple temporal patterns may require finer-grained segmentation. For example, for seasonal epidemics, time-series is smoother during the off-season and has more complex dynamics during outbreaks and times of higher incidence.
@@ 4 Related Works
Recent works have shown the efficacy of self-supervised representation learning for time-series for various classification and forecasting tasks in a wide range of applications such as modeling behavioral datasets Merrill and Althoff (2022); Chowdhury et al. (2022), power generation Zhang et al. (2019), health care Zhang et al. (2022). Franceschi et al. (2019) used triplet loss to discriminate segments of the same time-series from others. These works use methods such as contrastive losses Eldele et al. (2021); Yue et al. (2022) or other similarity metric-based techniques Tonekaboni et al. (2021). However, all these methods apply SSL on the same dataset that is used for training and may not adapt well to multi-domain datasets. There has been some recent works leveraging foundational models like LLMs for time-series forecasting across multiple applications. One set of works directly uses LLMs without fine-tuning to perform time-series forecasting via careful prompting Gruver et al. (2024); Jin et al. (2024); Liu et al. (2024d). Other works fine-tune LLMs specifically for time-series forecasting Zhou et al. (2023); Rasul et al. (2023); Ansari et al. (2024). The time-series representation used by these models includes using individual time-steps as input Ansari et al. (2024), converting each digit of the time-series to character embeddings to be directly used by LLMs Gruver et al. (2024); Jin et al. (2024) and uniform segmentation Das et al. (2023b); Woo et al. (2024); Zhou et al. (2023). LPTM performs superior to these methods while being 10x to 100x smaller than the large LLMs used as backbones.
@@ 5 Experiment Setup
We derive pre-train time-series datasets from multiple domains: Epidemics: We use a large number of epidemic time-series aggregated by Project Tycho (van Panhuis et al., 2018). from 1888 to 2021 for different diseases collected at state and city levels in the US. We remove time series with missing data and use time series for 11 diseases of very diverse epidemic dynamics such as seasonality, biology, geography, etc.: Hepatitis A, measles, mumps, pertussis, polio, rubella, smallpox, diphtheria, influenza, typhoid and Cryptosporidiosis (Crypto.). Electricity: We use ETT electricity datasets (ETT1 and ETT2) collected from (Zhou et al., 2021) at 1 hour intervals over 2 years. We use the default 12/4/4 train/val/test split and use the train split for pre-training as well. Traffic Datasets: We use 2 datasets related to traffic speed prediction. PEMS-Bays (PEM-B) and METR-LA (Li et al., 2017) are datasets of traffic speed at various spots collected by the Los Angeles Metropolitan Transportation Authority and California Transportation Agencies over 4-5 months. Demand Datasets: We use bike and taxi demand datasets (NY-B, NY-T) from New York City collected from April to June 2016 sampled every 30 minutes. We all but the last 5 days of data for training and pre-training. Stock forecasting: We also collect the time-series of daily stock prices of Nasdaq and S&P 500 index using Yahoo finance data (Xu and Berkely, 2014) from July 2014 to June 2019. We train and pre-train using the first 800 trading days and use the last 400 for testing. M4 competition time-series: We also used the 3003 time-series of M4 forecasting competition (Makridakis and Hibon, 2000) which contains time-series from multiple domains including demographics, finance, and macroeconomics. Motion and behavioral sensor datasets: We use the set of sensor datasets extracted from UEA archive (Bagnall et al., 2018) and UCI Machine learning repository (Asuncion and Newman, 2007) similar to (Chowdhury et al., 2022). Note that our publicly accessible pre-training dataset is significantly smaller than other pre-trained datasets used by past work Ansari et al. (2024); Das et al. (2023b) some of which use confidential data inaccessible to us. We also do not use any synthetic datasets like Das et al. (2023b); Ansari et al. (2024).
@@
We compared LPTM’s performance in various time-series tasks against twenty two state-of-the-art general forecasting and domain-specific baselines. First, we compare against recent pre-trained foundational time-series models: (1) LLM-Time Gruver et al. (2024), (2) TimesFM Das et al. (2023b), (3) Lag-LLAMA Rasul et al. (2023), (4) Chronos Ansari et al. (2024) and (5) MOIRAI Woo et al. (2024). We skip models like Time-LLM Jin et al. (2024), MOMENT Goswami et al. (2024) and Autotimes Liu et al. (2024b) which cannot perform zero-shot forecasting across domains and are outperformed by the aforementioned recent models when fine-tuned. We compared with (6) Informer Zhou et al. (2021), (7) Autoformer Chen et al. (2021), (8) iTransformer Liu et al. (2023) and (9) PatchTST Nie et al. (2022), four state-of-the-art transformer-based forecasting models. We also compare against other recent model (10) MICN (Wang et al., 2022), (11) TiDE Das et al. (2023a) (12) TFT Lim et al. (2021) and (13) TimesNeT Wu et al. (2023). We also compare with (14) N-HITS Challu et al. (2023) which uses multi-scale interpolation and (15) AutoARIMA Hyndman and Khandakar (2008) a ARIMA based model that does automatic hyperparameter search. We also compare it against three other state-of-art self-supervised methods for time-series: (16) TS2Vec (Yue et al., 2022), (17) TS-TCC (Eldele et al., 2021) and (18) SimMTM Dong et al. (2024) uses masking as pre-trained task for time-series classification.
@@ 6 Results
| Model | Flu-US | Flu-Japan | ETT1 | ETT2 | PEM-B | NY-B | NY-T | Nasdaq | M4 | Rank |
| AutoARIMA | 2.14 | 1344 | 0.73 | 0.64 | 4.1 | 4.13 | 16.43 | 0.62 | 1.89 | 25.06 |
| Informer | 1.62 | 1139 | 0.57 | 0.71 | 3.1 | 2.89 | 12.33 | 0.83 | 1.055 | 15.89 |
| Autoformer | 1.41 | 1227 | 0.72 | 0.82 | 2.7 | 2.73 | 12.71 | 0.19 | 0.887 | 13.67 |
@@
| TimesNet | 1.04 | 1194 | 0.56 | 0.62 | 3.9 | 2.83 | 11.82 | 0.19 | 1.055 | 12.11 |
| TFT | 1.21 | 1876 | 0.52 | 0.51 | 4.6 | 2.95 | 12.55 | 0.24 | 1.18 | 16.11 |
| iTransformer | 1.14 | 1256 | 0.57 | 0.59 | 4.3 | 2.83 | 13.16 | 0.29 | 1.125 | 17.5 |
| LLM-Time | 1.21 | 1319 | 0.52 | 0.49 | 3.9 | 3.7 | 12.11 | 0.21 | 1.064 | 14.05 |
| TimesFM | 1.32 | 1214 | 0.58 | 0.49 | 3.7 | 2.8 | 12.19 | 0.22 | 1.07 | 13.44 |
| Lag-LLAMA | 1.46 | 1416 | 0.61 | 0.57 | 3.9 | 2.9 | 13.43 | 0.28 | 1.33 | 20.16 |
| Chronos | 1.21 | 1228 | 0.59 | 0.52 | 3.7 | 3.1 | 12.82 | 0.27 | 1.04 | 15.55 |
| MOIRAI | 1.31 | 1336 | 0.62 | 0.55 | 3.9 | 3.5 | 13.71 | 0.24 | 1.21 | 19.22 |
@@
| EpiFNP | 0.52 | 872 | 0.81 | 1.25 | 4.1 | 2.98 | 12.11 | 0.28 | 1.281 | 16.77 |
| ColaGNN | 1.65 | 694 | 0.72 | 1.19 | 3.9 | 3.19 | 14.97 | 0.25 | 1.185 | 19.22 |
| TS2Vec | 1.85 | 905.9 | 0.99 | 1.74 | 3.5 | 3.11 | 13.48 | 0.94 | 1.344 | 21.94 |
| SimMTM | 1.31 | 1289 | 0.61 | 0.55 | 3.4 | 3.1 | 12.79 | 0.28 | 1.284 | 17.94 |
| TS-TCC | 1.94 | 1134.6 | 0.75 | 1.29 | 3.3 | 2.97 | 15.55 | 0.76 | 1.274 | 21 |
@@
| LPTM-NoSegment | 0.93 | 766 | 0.57 | 0.55 | 3.2 | 3.17 | 14.96 | 0.27 | 1.146 | 13.72 |
| LPTM-NoPreTrain | 0.96 | 827 | 0.46 | 0.57 | 3.7 | 2.66 | 12.43 | 0.25 | 1.271 | 11.66 |
| LPTM-NoLinProb | 0.92 | 885 | 0.43 | 0.53 | 3.1 | 2.49 | 12.17 | 0.19 | 1.032 | 6.55 |
| LPTM-OnlyRandMask | 0.87 | 895 | 0.51 | 0.52 | 2.8 | 2.42 | 12.36 | 0.21 | 1.076 | 8.0 |
| LPTM-OnlyLastMask | 0.79 | 773 | 0.44 | 0.48 | 2.7 | 2.31 | 12.04 | 0.19 | 0.916 | 3.77 |
The code for implementation of LPTM and datasets are provided at anonymized link333www.github.com/AdityaLab/Samay and hyperparameters are discussed in the Appendix. LPTM consists of 10 layers for the transformer and overall has about 100M parameters which 2x to over 10x smaller than other pre-trained time-series models.
@@
We also train the models with training data to evaluate fine-tuned performance. The forecasting performance are shown in Table 2. We evaluated LPTM against seventeen other forecasting baselines. LPTM is either the first or a close second best-performing model in all the benchmarks. LPTM generally outperforms general time-series forecasting methods as well as 2recent pre-trained models in all benchmarks despite its much lower parameter count (2-10x lower). Further, it is competitive or superior in performance to domain-specific methods designed specifically for the given domains (such as EpiFNP and STEP). LPTM beats the previous state-of-art domain-specific baselines in five of the benchmarks and comes second in four more. Finally, LPTM improves upon the state-of-art on electricity forecasting, traffic forecasting, and M4 datasets.
@@
Unlike other autoregressive foundational models designed for forecasting, LPTM can be used for classification due to its encoder-style architecture. We add a single classification layer over pooled output embeddings to predict the class logits. We evaluate LPTM and baselines on the classification of 35 sensor and behavioral datasets from UCI classification repositiry(Asuncion and Newman, 2007). We report the accuracy scores in Table 3. We observe that LPTM has highest mean rank and largest number of times it outperforms all baselines.
| Informer | Autoformer | TimesNet | TARNet | TS2Vec | TS-TCC | TST | SimMTM | CRT | LPTM | |